Test of DX LGSW with CalUnit and rigid secondary

Hexapod settings with rigid secondary:

Right PSF 2ndary
X Y Z Rx Ry Rz
-4.75 -8.30 -5.15 4.97 -230.18 8
Right PSF tertiary
Tip Tilt Piston Rz
100 0 0 0

DX and SX LM2 adjustment

LM2 have been adjusted manually. For DX the nominal LM1 position was good. For SX LM1 had to be moved.
SX LM1 nominal positions
A1 18630
A2 22746
A3 15958

LgswUnitSx test with FLAO as TT sensor

The system is setup for FLAO BIN2 using parameters of yesterday, the light is well in place.

BOARD SETUP: 12 Dec 15 - ARGOS_BIN2

Center of FLAO FoV is at stages position: -77.5; -49.1mm. To pickup the white light source when controlling the system with the AARB an offset of +5.5; +2.4mm must be sent when lauching.

CalUnit white light source is switched on automatically by the AARB, when fully opened it is equivalent to a magnitude 9.0. This value must be set in the AARB GUI before launching.

TT loop closed on PYR BIN2; 150modes on LGSW; no disturb; JCL:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes Modalplot
20151212_164334 0.2,0.1,0.05 43.7437 20151211_202800 -1 T F Z @ 50mm 20151212_164334-20151212_164833_modalplot.png
20151212_164439 0.2,0.1,0.05 39.3759 20151211_202800 -1 T F
20151212_164624 0.2,0.1,0.05 117.293 20151211_202800 -1 T F Z @ 35mm
20151212_164658 0.2,0.1,0.05 129.95 20151211_202800 -1 T F
20151212_164814 0.2,0.1,0.05 118.03 20151211_202800 -1 T T
20151212_164833 0.2,0.1,0.05 112.584 20151211_202800 -1 T T

Then an 0.8" equivalent disturb is loaded, the loop is closed again, always keeping PCL even when AO loop was open:
TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes Modalplot
20151212_165457 1,0.45,0.45 74.6697 20151211_202800 -1 T T 20151212_165457-20151212_170432_modalplot.png
20151212_165522 1,0.45,0.45 73.2269 20151211_202800 -1 T T
20151212_165548 1,0.45,0.45 74.2999 20151211_202800 -1 T T
20151212_170400 1,0.45,0.45 72.921 20151211_202800 -1 T T
20151212_170421 1,0.45,0.45 73.1414 20151211_202800 -1 T T
20151212_170432 1,0.45,0.45 75.2314 20151211_202800 -1 T T
10:30: the BCU continues to stack, we fear a possible under-powering problem. We try to run the BCU at 500Hz but the problem persists...

We find out that when the FLAO slopes are enabled and the jitter loop is closed, the piezos start to oscillate. It seems that we are writing the FLAO slopes in the wrong BCU memory address...

LgswUnitSx AO calibrations

11:15: we switch to check LgswUnitSx calibrations: system setup at 500Hz. TCoG48 subaps are loaded (20151211_181800).

We close the loop on LGSW only and we acquire the best flat for the sensor.

ADSEC FLAT: 20151212_182300_ARGOS_150modes

HEXAPOD: 4.57, 0.05, 7.92, 228, -115

Beam X Y
Blue 325 421
Yellow 376 438
Red 487 432

# modes IM REC IM LGSW REC
150 0 20151212_182939 red pupil has ear on SY tip
20151212_183322 red motor dowm by 20 steps, less ear
20151212_183634 down by other 20 steps, lower but maybe caused by lower signals
20151211_183209_IM.png

Beam X Y
Blue 320 431
Yellow 376 448
Red 495 432
Jitter initial position: 20151212_184142

Pupil target tag: 20151212_184534

# modes IM REC IM LGSW REC
150 0 20151212_184343 20151212_184500
We are running at 500Hz and LB complains, so we set 610Hz and we repeat the IM:

# modes IM REC IM LGSW REC
150 0 20151212_185051 -
150 0 20151212_185358 -
On both on them the red pupil shows ear: we discover that the jitter loop was closed.

# modes IM REC IM Notes
150 0 20151212_185725 600Hz, JOL
150 0 20151212_190109 370Hz, JOL
150 0 20151212_191843 530Hz, JOL
PBL: we find that the argos.lgsw.controllergetFrameRateFromControllerInHz is oscillating between 400 and 800Hz while BCU reads just 570Hz...

20151212_193200_getFrameRateFromController.png

we try to generate a modal history having 10 steps up and 10 steps down per mode (that's actually no push pull) to wash out the frame rate oscillation.

Modal amplitudes Modal history
20151211_162400_rms0.05 20151212_164600_150modes

# modes IM REC IM Notes
150 0 20151212_195337 600Hz, JCL
150 0 20151212_195812 600Hz, JOL
Last one seems the best one so far, but blue pupil is quite noisy... We try to offset the hexapod pointing by [0.0; 0.2]arcsec.

# modes IM REC IM LGSW REC Notes
150 0 20151212_200725 20151212_200900 worse
150 20151212_200900 20151212_201153 not good
We change again the hexapod pointing by [0.0; -0.4]arcsec. Jitter initial position: 20151212_201607.

# modes IM REC IM LGSW REC Notes
150 20151212_200900 20151212_201810 20151212_202000 -
150 20151212_202000 20151212_202229 20151212_203300 very good, TT removed
150 20151212_220600 w/o TT removal
We try again to acquire an IM with push-pull: 3 steps up and 3 down:

# modes IM REC IM LGSW REC Notes
150 20151212_203300 20151212_205038 - bad
We try to disable external clock and to run on the LGSW internal one to see if the frame rate stops to oscillate.

# modes IM REC IM LGSW REC Notes
150 20151212_203300 20151212_205258 - bad ear
We fear a problem with OVS.

14:30: we also fear that maybe the TT amplitudes are too small, so we use the original DX modal amplitudes to biuld the modal history: 20151212_213400_150modes.

# modes IM REC IM LGSW REC Notes
150 0 20151212_213952 - better Tip Tilt signal, x5 wrt before
150 0 20151212_214224 20151212_215700  
We build a combo rec on last push-pull IM and also on the 10 steps ones:

REC TAG LGSW REC COMBO REC
TCoG48 only w/ push-pull 20151212_215700 20151212_215800
TCoG48 only 10 steps 20151212_203300 20151212_214900
TCoG48 only 10 steps, w/o TT rem 20151212_220600 20151212_220700
We test these Combo recs, w/o disturb they oscillates between 0.4 and 0.5. We load 1.0" disturb:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Modalplot
20151212_221029 0.5,0.5,0.5 154.264 20151212_220700 -1 F F 20151212_221029-20151212_221050_modalplot.png
20151212_221050 0.5,0.5,0.5 142.308 20151212_220700 -1 F F
20151212_221432 0.5,0.5,0.5 945.461 20151211_183400 -1 F F 20151212_221432-20151212_221722_modalplot.png
20151212_221458 0.4,0.4,0.4 188.554 20151211_183400 -1 F F
20151212_221722 0.45,0.45,0.45 423.74 20151211_183400 -1 F F
20151212_222305 0.45,0.45,0.45 162.393 20151212_220700 -1 F F 20151212_224603-20151212_224648_modalplot.png
20151212_224603 0.2,0.1,0.05 327.924 20151212_220700 -1 F F
20151212_224648 0.2,0.1,0.05 341.967 20151212_220700 -1 F F
W/o disturb:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Modalplot
20151212_225001 0.2,0.1,0.05 215.681 20151212_220700 -1 F F 20151212_225001-20151212_225020_modalplot.png
20151212_225020 0.2,0.1,0.05 228.098 20151212_220700 -1 F F
Best Z stage position for N30 camera is 52mm.

We combine now LGSW recs with FLAO ones:

REC TAG LGSW REC FLAO REC COMBO REC
TCoG48 10 steps w/o TT rem, BIN2 20151212_220600 20151202_213530 20151212_233400
TCoG48 10 steps w/o TT rem, BIN4 20151212_220600 20151204_191745 20151212_233600
20:00: CalUnit barrel was shifted inside the gusset to work with hexapod at 0mm in Z.

HEXA 4.37 -0.21 -0.8 274 -141
Jitter initial position: 20151213_033245

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Modalplot
20151213_033548 0.45,0.45,0.45 126.481 20151212_220700 -1 F F 20151213_033548-20151213_033619_modalplot.png
20151213_033619 0.45,0.45,0.45 117.352 20151212_220700 -1 F F
Board @: -72.6; -47.3; 52.0 mm

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Modalplot
20151213_040357 0.45,0.45,0.45 158.845 20151212_220700 -1 F F 20151213_040357-20151213_040437_modalplot.png
20151213_040437 0.45,0.45,0.45 158.313 20151212_220700 -1 F F

SX quad-cell calibrations

21:30: we switch on the QC. LUCI rotator set to 0deg. Pupil rerotator 258.9deg. Pyramid with BIN2.

  • Board @: -74.164; -45.865; 52.0 mm -> to center Pyramid

  • Board @: -73.95; -46.6; 52.0 mm -> to center QC

Pyramid to TT mirror IM: 20151213_052700

  • Board @: -74.65; -51.45 -> to center stages on pennina
  • FW2 @ -0.015 filters

Quad-cell IM: 20151213_062235

Quad-cell IM: 20151213_063244 -> QC REC: 20151213_063500 acquired with 1" defocus

We combine the recs:

REC TAG LGSW REC QC REC COMBO REC
TCoG48 10 steps w/o TT rem, QC 20151212_220600 20151213_063500 20151213_063700
And we test them. To center quad cell we must send an offset of [3.3; 3.3]mm from the AArb.

Hot pixel on CCD47 @ BIN1: [547;624] with 90-10 + empty filters, [518;648] with 600-1000 + empty, [487;602] with mirror + empty.

Board positions for QC: -74.4; -46.0 -> offset of [3.1;3.1]mm

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Modalplot
20151213_080915 1.5,0.5,0.5 70.3386 20151213_063700 -1 T F 20151213_080915-20151213_081626_modalplot.png
20151213_080931 1.5,0.5,0.5 76.2197 20151213_063700 -1 T F
20151213_081043 1.5,0.4,0.4 80.6916 20151213_063700 -1 T F
20151213_081057 1.5,0.4,0.4 83.0237 20151213_063700 -1 T F
20151213_081626 4,0.45,0.45 89.3721 20151213_063700 -1 T F
Then we build a QC only combined rec:

REC TAG LGSW REC QC REC COMBO REC
QC only zeros 20151213_063500 20151213_020000
-- MarcoBonaglia - 12 Dec 2015
Topic revision: r5 - 08 Jan 2016, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback